OpenStudy (anonymous):

does any one know what 3x+4y=6 5x+2y=0

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

3x+4y=6 y=-3/4 (x-2)

OpenStudy (anonymous):

3x+4y=6 - 2(5x+2y)=0 ------------ -7x=6